# Sparse-IFT: Sparse Iso-FLOP Transformations for Maximizing Training Efficiency

### v3

@@ Appendix B Graph Analysis of Sparse-IFT with DST
Figure 6: Illustrates the dynamic interplay between the (top row) Iterative Mean Difference Bound, $\Delta r_{imdb}$ and test accuracy, as well as the correlation between the Ramanujan Gap, $\Delta r$ and test accuracy throughout the training process. This illustrates the evolving relationship between spectral graph properties and network performance, shedding light on the connectivity dynamics of the Sparse-IFT networks trained with DST.
@@ C.1 Computer Vision: Pre-Training Settings
Our implementation of ImageNet follows the standard setup from (Krizhevsky et al., 2017; Simonyan & Zisserman, 2014). The image is resized with its shorter side randomly sampled in [256, 480] for scale augmentation (Simonyan & Zisserman, 2014). A 224 $\times$ 224 crop is randomly sampled from an image or its horizontal flop, and then normalized. For evaluation, the image is first resized to 256 $\times$ 256, followed by a 224 $\times$ 224 center crop, and then normalized. Following recent advances in improving ResNets, we initialize the network with Kaiming He initialization (He et al., 2016) and zero-init residuals (He et al., 2019).
@@ D.1 Details for GPT End-to-End Training
We demonstrate the benefits of using Sparse-IFT transformations in the NLP domain by pre-training GPT-3 models and performing zero-shot eval on downstream tasks from the HuggingFace Open LLM leaderboard. Here, we pre-train the models on the Pile (Gao et al., 2020) dataset. To train all GPT models, we use the AdamW optimizer (Loshchilov & Hutter, 2017) with $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$. The global norm is clipped at 1.0, and a weight decay of 0.1 is used. There is a learning rate warmup over the first 375M tokens, followed by a cosine decay to 10% of the peak learning rate. We follow the recently published Chinchilla (Hoffmann et al., 2022) recommendations for obtaining loss-optimal pre-trained baseline configurations of models. The context window size is 2048 following (Brown et al., 2020). Table 13 shows a detailed breakdown of the model architectures, learning rate, and training settings.
In Tables 13 and 14, we outline the architecture configurations for Sparse Wide IFT 50% and 75% variants. We train the Sparse Wide GPT-3 models using the dynamic sparse training algorithm, SET (Mocanu et al., 2018) on the Cerebras CS-2 to realize the acceleration from unstructured sparsity. Currently, Cerebras CS-2’s specialized kernels support training with dynamic unstructured sparsity via SET; therefore, results in this section are reported with SET. In Tables 13 and 14, $n_{params}$ is the total number of trainable parameters, $n_{layers}$ is the number of decoder layers, and $d_{model}$ is the base size of the model. The feedforward bottleneck is four times the base size, i.e., $d_{\text{ff}}=4\times d_{model}$. Finally, $n_{heads}$ is the number of attention heads, and $d_{head}$ is the dimension of each attention head.
@@
In this section we provide additional details on the benchmarking setups for inference on Neural Magic DeepSparse (NeuralMagic, 2021; Iofinova et al., 2021; Kurtz et al., 2020) sparsity-aware runtime and training on the Cerebras CS-2 (Lie, 2022; Cerebras, 2023) for evaluating the efficiency of Sparse-IFT with respect to the wall-clock time.
@@
The results in Figure 5 of Section 6 show that up 75%, there is minimal computational overhead compared to training the original dense baseline model. At 90% sparsity, our results lie between these two spectrums (green line). Using Neural Magic’s sparse inference runtime, we observe a significant reduction in inference latency, bringing down the relative increase in latency from 19.5x to 3.5x. Similiarly, in the case of training on the Cerebras CS-2, we observe a significant reduction in training-time, bringing down the relative increase from 10.6x to 2.8x.
@@
Appendix F Broader Impact and Ethical Considerations
The landscape of machine learning (ML) has witnessed an exponential growth in models, particularly in domains such as natural language processing and computer vision. However, this surge in model size has come with a considerable cost in terms of compute, memory, and energy requirements. Our approach, Sparse Iso-FLOP Transformations (Sparse-IFT), represents a significant stride toward mitigating these resource-intensive demands.
Sparse-IFT introduces a novel approach that enhances the efficiency of training large neural networks. Remarkably, it achieves improved accuracy while maintaining the same FLOPs as the original dense baseline model. Our method holds promise for positive environmental impacts, given the substantial computational resources typically associated with training large neural networks.
The potential sustainability contribution lies in the fact that, as sparse ML software and hardware co-design continues to evolve, we may be able to train more accurate ”larger sparse” networks within the confines of the same computational budget as a smaller dense model. This paradigm shift could usher in a more environmentally conscious approach to deep learning, addressing the concerns associated with the escalating resource requirements of ever-expanding models.
The key lies in achieving sparsity acceleration, which necessitates a harmonious collaboration between hardware support for sparsity and the development of sparse ML techniques. The seamless integration of these elements is crucial for realizing the full potential of Sparse-IFT and similar innovations. This holistic co-design approach ensures that hardware architectures are optimized to complement sparse techniques, fostering a sustainable and efficient trajectory for the future of deep learning.
As we continue to explore the intersection of hardware support for sparsity and the evolution of sparse ML techniques, our benchmarking analysis in Section 6 and Appendix E serve as a practical illustration of the transformative potential of Sparse-IFT. It not only substantiates the theoretical promises but also offers a roadmap for future developments in the pursuit of sustainable and efficient deep learning practices.
Appendix G Author Contributions
@@
  • Vithursan Thangarasa was an integral part of the project by participating in discussions with Shreyas Saxena and contributing to the method. He also implemented all Sparse-IFT transformations in PyTorch, proposed using non-linearity in Sparse-IFT, conducted experiments for the entire study on CIFAR-100 and its ablations, obtained initial results on ImageNet, extended Sparse-IFT to efficient architectures (e.g., BotNet, MobileViT), conducted the entire study with GPT on Cerebras CS-2, and contributed to writing parts of the manuscript.
  • Shreyas Saxena conceived the key idea of matching the FLOPs of Sparse Wide transformation to a compact dense model, extended the idea to other members of the Sparse-IFT family, helped with the implementation, established cardinality of Sparse-IFT members to explain the results, conducted experiments for BERT, benchmarked Sparse-IFT for inference, and wrote majority of the manuscript.
  • Abhay Gupta validated sparse optimizers in PyTorch, conducted experiments with Sparse-IFT ResNet variants on ImageNet, obtained results with MobileNetV2 architecture, helped with pre-training of Sparse-IFT variants of GPT on Cerebras CS-2, conducted all experiments of Sparse-IFT on downstream CV tasks, and contributed to writing parts of the manuscript.
  • Sean helped with the bring-up of sparsity support on Cerebras CS-2 which was crucial for benchmarking and training Sparse-IFT variants of GPT models, and provided feedback to improve the structuring and presentation of the manuscript.

### v4

@@ Appendix B Graph Analysis of Sparse-IFT with DST
Figure 6: Top row illustrates the dynamic interplay between the Iterative Mean Difference Bound, $\Delta r_{imdb}$ and test accuracy, and bottom row shows the correlation between the Ramanujan Gap, $\Delta r$ and test accuracy throughout the training process. This illustrates the evolving relationship between spectral graph properties and network performance, shedding light on the connectivity dynamics of the Sparse-IFT networks trained with DST.
@@ C.1 Computer Vision: Pre-Training Settings
Our implementation of ImageNet follows the standard setup from (Krizhevsky et al., 2012; Simonyan & Zisserman, 2014). The image is resized with its shorter side randomly sampled in [256, 480] for scale augmentation (Simonyan & Zisserman, 2014). A 224 $\times$ 224 crop is randomly sampled from an image or its horizontal flop, and then normalized. For evaluation, the image is first resized to 256 $\times$ 256, followed by a 224 $\times$ 224 center crop, and then normalized. Following recent advances in improving ResNets, we initialize the network with Kaiming He initialization (He et al., 2016) and zero-init residuals (He et al., 2019).
@@ D.1 Details for GPT End-to-End Training
We demonstrate the benefits of using Sparse-IFT transformations in the NLP domain by pre-training GPT-3 models and performing zero-shot eval on downstream tasks from the HuggingFace Open LLM leaderboard. Here, we pre-train the models on the Pile (Gao et al., 2020) dataset. To train all GPT models, we use the AdamW optimizer (Loshchilov & Hutter, 2017) with $\beta_{1}=0.9$, $\beta_{2}=0.95$ and $\epsilon=10^{-8}$. The global norm is clipped at 1.0, and a weight decay of 0.1 is used. There is a learning rate warmup over the first 375M tokens, followed by a cosine decay to 10% of the peak learning rate. We follow the recently published Chinchilla (Hoffmann et al., 2022) recommendations for obtaining loss-optimal pre-trained baseline configurations of models. The context window size is 2048 following (Brown et al., 2020). Table 13 shows a detailed breakdown of the model architectures, learning rate, and training settings.
In Tables 13 and 14, we outline the architecture configurations for Sparse Wide IFT 50% and 75% variants. We train the Sparse Wide GPT-3 models using the dynamic sparse training algorithm, SET (Mocanu et al., 2018) on the Cerebras CS-2 to realize the acceleration from unstructured sparsity. Currently, Cerebras CS-2’s specialized kernels support training with dynamic unstructured sparsity via SET; therefore, results in this section are reported with SET. In Table 13, $n_{params}$ is the total number of trainable parameters, $n_{layers}$ is the number of decoder layers, and $d_{model}$ is the base size of the model. The feedforward bottleneck is four times the base size, i.e., $d_{\text{ff}}=4\times d_{model}$. Finally, $n_{heads}$ is the number of attention heads, and $d_{head}$ is the dimension of each attention head.
@@
In this section we provide additional details on the benchmarking setups for inference on Neural Magic DeepSparse (NeuralMagic, 2021; Iofinova et al., 2021; Kurtz et al., 2020) sparsity-aware runtime and training on the Cerebras CS-2 (Lie, 2023; Cerebras, 2023) for evaluating the efficiency of Sparse-IFT with respect to the wall-clock time.
@@
The results in Figure 5 of Section 6 show that up 75%, there is minimal computational overhead compared to training the original dense baseline model. At 90% sparsity, our results lie between these two spectrums (green line). Using Neural Magic’s sparse inference runtime, we observe a significant reduction in inference latency, bringing down the relative increase in latency from 19.5x to 3.5x. Similiarly, in the case of training on the Cerebras CS-2, we observe a significant reduction in training-time, bringing down the relative increase from 10.6x to 2.8x. The dense GPT-3 model achieved a throughput of 828.71 iterations per second on the CS-2, while the Sparse Wide IFT variants recorded throughputs of 637.5, 595.3, and 294.3 at respective sparsity levels, resulting in overheads of 1.30x, 1.39x, and 2.82x, respectively.
@@
Appendix F Author Contributions
@@
  • Vithursan Thangarasa was an integral part of the project by participating in discussions with Shreyas Saxena and contributing to the method. He also implemented all Sparse-IFT transformations in PyTorch, proposed using non-linearity in Sparse-IFT, analyzed DST methods via spectral analysis, conducted experiments for the entire study on CIFAR-100 and its ablations, obtained initial results on ImageNet, extended Sparse-IFT to efficient architectures (e.g., BotNet, MobileViT), and contributed to writing several sections of the manuscript.
  • Shreyas Saxena conceived the key idea of matching the FLOPs of Sparse Wide transformation to a compact dense model, extended the idea to other members of the Sparse-IFT family, helped with the implementation, established cardinality of Sparse-IFT members to explain the results, benchmarked Sparse-IFT for inference, and wrote several sections of the manuscript.
  • Abhay Gupta validated sparse optimizers in PyTorch, conducted experiments with Sparse-IFT ResNet variants on ImageNet, helped with pre-training of Sparse-IFT variants of GPT on Cerebras CS-2, conducted all experiments of Sparse-IFT on downstream CV tasks, and contributed to writing parts of the manuscript.
  • Sean Lie helped with the bring-up of sparsity support on Cerebras CS-2 which was crucial for benchmarking and training Sparse-IFT variants of GPT models, and provided feedback to improve the structuring and presentation of the manuscript.